OpenStudy (anonymous):

Determine whether the system of linear equations has one and only one solution, infinitely many solutions, or no solution. 2x − y = 5 3x + y = −7

OpenStudy (anonymous):

so far i hve 5x=-2? this is right

OpenStudy (anonymous):

Yup, you got it right so far.

OpenStudy (anonymous):

so, do i hve to divide -2 by 5 now?

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

okay, so i got.-0.4

OpenStudy (anonymous):

Perfect, now substitute it into either equation to come up with the value of the other variable.

OpenStudy (anonymous):

okay so now i have 2(-0.4)-y=5. den -0.8-y=5. dont i add 0.8 to 5? if so i got 4.2? i den plug it into one of the equations 2x-y=5 and the answer isnt the same. wat did i do wrong?